a small block b is placed on a block a of mass 5kg and lenght 20cm.initially block b is near the right end of block a.


a constant horizontal force of 10newton is applied on block a


all the surfaces are assumed frictionless


find the time elapsed before the block b seperates from block a


 


 


 


 


plz help me

Grade:10

2 Answers

vikas askiitian expert
13 years ago

let the mass of block b is mb....

force on lower block is F=10N....

accleration of system is a...

    (mb+5)a = 10

         a= 10/(mb+5) m/s2

now a seudo force acts on the upper block in opposite direction and its magnitude is f=(mb)a...

 due to this force block moves in opposite direction with accleration a1....

   mba1=mba

      a1=a=10/(5+mb)

when the block covers a distance of 20cm then it separates out...

   0.2 = a1t2/2

   t2  =  (0.4/a1)=[0.04(5+mb)]

   t=  0.2(5+mb)1/2

Sayan
6 years ago
The block A will accelerate under the force 10 N applied on it. But the surfaces being frictionless, block B will remain stationary. Soblock B will separate from block A when block A has travelled a distance s equal to its length i.e. 20cm.For block A:u=0 m/s; a=Force/Mass=10/5=2NSo, v²=√(u²+2as)v=√2as=√(2*2*0.2)=0.894 m/sTherefore, v=u+at0.894=0+2t=>t=0.447 sec=0.45 sec